# Grandmaster-Level Chess Without Search

### v1

@@ Acknowledgments
We thank Aurélien Pomini, Avraham Ruderman, Eric Malmi, Charlie Beattie, Chris Colen, Chris Wolff, David Budden, Dashiell Shaw, Guillaume Desjardins, Hamdanil Rasyid, Himanshu Raj, Joel Veness, John Schultz, Julian Schrittwieser, Laurent Orseau, Lisa Schut, Marc Lanctot, Marcus Hutter, Matthew Aitchison, Nando de Freitas, Nenad Tomasev, Nicholas Carlini, Nick Birnie, Nikolas De Giorgis, Ritvars Reimanis, Satinder Baveja, Thomas Fulmer, Tor Lattimore, Vincent Tjeng, Vivek Veeriah, and Zhengdong Wang for insightful discussions and their helpful feedback.
@@ Appendix A Experimental Setup
State-Value | Behavioral Cloning | Action-Value | |||||
Split | Games | Records | Bytes | Records | Bytes | Records | Bytes |
Train | 591 897 | 43.7 MB | 589 130 | 41.1 MB | 17 373 887 | 1.4 GB | |
5 747 753 | 422.0 MB | 5 720 672 | 397.4 MB | 167 912 926 | 13.5 GB | ||
55 259 971 | 4.0 GB | 54 991 050 | 3.8 GB | 1 606 372 407 | 129.0 GB | ||
530 310 443 | 38.6 GB | 527 633 465 | 36.3 GB | 15 316 914 724 | 1.2 TB | ||
Test | 62 829 | 4.6 MB | 62 561 | 4.4 MB | 1 838 218 | 148.3 MB |
@@ A.2 Main Setup
Concretely our base setup is as follows: We train for 20 million steps with a batch size of 4096, meaning that we train for 5.35 epochs. We use the Adam optimizer (Kingma and Ba, 2015) with a learning rate of 1e-4. We train on the dataset generated from 10 million games (cf. Table A1) for the action value policy with 128 return buckets and a stockfish time limit of 0.05s. We use the unique sampler and Polyak averaging for evaluation and evaluate on 1000 games (cf. Table A1) and 1000 puzzles from a different month than that used for training.
We train a vanilla decoder-only transformer without causal masking (Vaswani et al., 2017), with the improvements proposed in LLaMA (Touvron et al., 2023a, b), i.e., post-normalization and SwiGLU (Shazeer, 2020). We use three different model configurations: (i) 8 heads, 8 layers, and an embedding dimension of 256, (ii) 8 heads, 8 layers, and an embedding dimension of 1024, and (iii) 8 heads, 16 layers, and an embedding dimension of 1024.
@@ A.3 Ablation Setup
Concretely our base setup is as follows: We train for 5 million steps with a batch size of 1024, meaning that we train for 3.19 epochs. We use the Adam optimizer (Kingma and Ba, 2015) with a learning rate of 4e-4. We train on the dataset generated from 1 million games (cf. Table A1) for the action value policy with 32 return buckets and a stockfish time limit of 0.5s. We use the unique sampler and train a vanilla decoder-only transformer (Vaswani et al., 2017) with post-normalization, 8 heads, 8 layers, an embedding dimension of 256, and no causal masking. We use Polyak averaging for evaluation and evaluate on 1000 games (cf. Table A1) and 1000 puzzles from a different month than that used for training.
@@ A.4 Dataset Statistics
We visualize some dataset statistics in Figures A1 and A2.
A.5 Playing-strength evaluation on Lichess
@@
We evaluate and calibrate the playing strength of our models by playing against humans and bots on Lichess. Our standard evaluation allows for both playing against bots and humans (see Table 1), but since humans tend to rarely play against bots the Elo ratings in this case are dominated by playing against other bots (see our discussion of how this essentially creates two different, somewhat miscalibrated, player pools in Section 5). In our case the policies in the column denoted with ‘vs. Bots’ in Table 1 have played against some humans but the number of games against humans is of total games played. To get better calibration against humans we let our largest model play exclusively against humans (by not accepting games with other bots) which leads to a significantly higher Elo ranking (see Table 1). Overall we have played the following numbers of games for the different policies shown in Table 1: M (553 games), M (169 games), M (228 games against bots, 174 games against humans), Stockfish (30 games), GPT-3.5-turbo-instruct (181 games).
A.6 Stockfish and AlphaZero Setup
Stockfish
@@ A.6 Stockfish and AlphaZero Setup
We use Stockfish 16 (the version from December 2023) throughout the paper. When we play, we use the oracle we used for training, which is an unconventional way to play with this engine: We evaluate each legal move in the position for 50ms, and return the best move based on these scores. This is not entirely equivalent to a standard thinking time of 50ms times the number of legal moves per position, as we force Stockfish to spend 50ms on moves that could be uninteresting and unexplored. We chose to keep this setup to have a comparison to the oracle we train on. Note that, when comparing the legal moves in a given position, we do not clear Stockfish’s cache between the moves. Therefore, due to the way the cache works, this biases the accuracy of Stockfish’s evaluation to be weaker for the first moves considered. Finally, due to the nature of our internal hardware setup, we use two different kinds of chips to run Stockfish: (i) to compute the Lichess Elo, we use a 6-core Intel(R) Xeon(R) W-2135 CPU @ 3.70GHz, and (ii) to compute the tournament Elo, we use a single Tensor Processing Unit (V3), as for all the other agents.
AlphaZero
@@
We use the AlphaZero version from 2020, with a network trained at that time (Schrittwieser et al., 2020). We use three different versions: (i) policy network only, (ii) value network only and (iii) standard version with search. For (i), we are using the probability distribution over actions returned by the policy network, and take the argmax as the best action. For (ii), we do a search limited to depth 1, with 100 MCTS simulations, enough to cover all legal actions, and take the argmax over visit counts. For (iii), we use the standard search from the paper, with 400 MCTS simulations and the exact same UCB scaling parameters. We also take the argmax over visit counts. Note that AlphaZero’s policy and value network have been trained on M games, whereas we trained our largest models on only M games.
A.7 Computational Resources
@@
Our codebase is based on JAX (Bradbury et al., 2018) and the DeepMind JAX Ecosystem (DeepMind et al., 2020; Hennigan et al., 2020). We used 4 Tensor Processing Units (V5) per model for the ablation experiments. We used 128 Tensor Processing Units (V5) per model to train our large (9M, 136M and 270M) models. We used a single Tensor Processing Unit (V3) per agent for our Elo tournament.
@@ Appendix B Additional Results
B.1 Loss Curves
@@
In Figure A4 we visualize the train and test loss curves for the scaling experiment from Section 3.3. In line with the results shown in the main paper we observe that on the smallest training set, models with M parameters start to overfit, but not for the larger training sets. Except for the overfitting cases we observe that larger models improve both the training and test loss, regardless of training set size, and that larger training set size improves the test loss when keeping the model size constant.
B.2 Predictor-Target Comparison
@@
Throughout all these results we observe lower performance of the behavioral cloning policy, despite being trained on a comparable number of datapoints as the state-value policy. The main hypothesis for this is that the amount of information in the behavioral cloning dataset is lower than the state value dataset, since we throw away any information in the state- or action-values beyond the index of the oracle action. We suspect that training on the full action distribution of the oracle (with cross-entropy loss), rather than the best action only would largely close this gap, but we consider this question beyond the scope of this paper and limit ourselves to simply reporting the observed effect in our setting.
B.3 Polyak Averaging
We investigate the impact of Polyak averaging, an optimization technique where parameters are set to a weighted average over the last iterations rather than just using the most recent value, using the same setup as for our ablation experiments (see Section A.3). When using Polyak averaging with an exponential moving average decay factor of 0.99, we obtain a Kendall’s of 0.259, a puzzle accuracy of 83.3%, and an action accuracy of 63.0%. In contrast, standard evaluation, obtains a Kendall’s of 0.258, a puzzle accuracy of 83.1%, and an action accuracy of 62.8%. Thus, we use Polyak averaging for all experiments.
B.4 Tactics
@@
B.5 Playing Style
@@
We recruited chess players of National Master level and above to analyze our agent’s games against bots and humans on the Lichess platform. They made the following qualitative assessments of its playing style and highlighted specific examples (see Figure A8). Our agent has an aggressive enterprising style where it frequently sacrifices material for long-term strategic gain. The agent plays optimistically: it prefers moves that give opponents difficult decisions to make even if they are not always objectively correct. It values king safety highly in that it only reluctantly exposes its own king to danger but also frequently sacrifices material and time to expose the opponent’s king. For example 17 .. Bg5 in game B.5.1 encouraged its opponent to weaken their king position. Its style incorporates strategic motifs employed by the most recent neural engines (Silver et al., 2017; Sadler and Regan, 2019). For example it pushes wing pawns in the middlegame when conditions permit (see game B.5.2). In game B.5.3 our agent executes a correct long-term exchange sacrifice. In game B.5.4 the bot uses a motif of a pin on the back rank to justify a pawn sacrifice for long term pressure. Game B.5.5 features a piece sacrifice to expose its opponent’s king. The sacrifice is not justified according to Stockfish although the opponent does not manage to tread the fine line to a permanent advantage and blunders six moves later with Bg7.
@@
Our agent has a distinct playing style to Stockfish: one analyzer commented “it feels more enjoyable than playing a normal engine”, “as if you are not just hopelessly crushed”. Indeed it does frequently agree with Stockfish’s move choices suggesting that our agent’s action-value predictions match Stockfish’s. However the disagreements can be telling: the piece sacrifice in the preceding paragraph is such an example. Also, game B.5.6 is interesting because our agent makes moves that Stockfish strongly disagrees with. In particular our agent strongly favours 18 .. Rxb4 and believes black is better, in contrast Stockfish believes white is better and prefers Nd4. Subsequent analysis by the masters suggests Stockfish is objectively correct in this instance. Indeed on the very next move our agent has reversed its opinion and agrees with Stockfish.
@@
Our agent’s aggressive style is highly successful against human opponents and achieves a grandmaster-level Lichess Elo of 2895. However, we ran another instance of the bot and allowed other engines to play it. Its estimated Elo was far lower, i.e., 2299. Its aggressive playing style does not work as well against engines that are adept at tactical calculations, particularly when there is a tactical refutation to a sub-optimal move. Most losses against bots can be explained by just one tactical blunder in the game that the opponent refutes. For example Bxh3 in game B.5.7 loses a piece to g4.
@@
Finally, the recruited chess masters commented that our agent’s style makes it very useful for opening repertoire preparation. It is no longer feasible to surprise human opponents with opening novelties as all the best moves have been heavily over-analyzed. Modern opening preparation amongst professional chess players now focuses on discovering sub-optimal moves that pose difficult problems for opponents. This aligns extremely well with our agent’s aggressive, enterprising playing style which does not always respect objective evaluations of positions.
B.5.1 King weakening game

### v2

@@ Acknowledgments
We thank Aurélien Pomini, Avraham Ruderman, Eric Malmi, Charlie Beattie, Chris Colen, Chris Wolff, David Budden, Dashiell Shaw, Guillaume Desjardins, Guy Leroy, Hamdanil Rasyid, Himanshu Raj, John Schultz, Julian Schrittwieser, Laurent Orseau, Lisa Schut, Marc Lanctot, Marcus Hutter, Matthew Aitchison, Matthew Sadler, Nando de Freitas, Neel Nanda, Nenad Tomasev, Nicholas Carlini, Nick Birnie, Nikolas De Giorgis, Ritvars Reimanis, Satinder Baveja, Thomas Fulmer, Tor Lattimore, Vincent Tjeng, Vivek Veeriah, Zhengdong Wang, and the anonymous reviewers for insightful discussions and their helpful feedback.
@@ A.2 Main Setup
We train for 10 million steps with a batch size of 4096, meaning that we train for 2.67 epochs on the full training dataset (M games). We use the Adam optimizer [17] with a learning rate of . We train on the dataset generated from 10 million games (cf. Table A1) for the action-value policy with 128 return buckets and a stockfish time limit of 0.05s. We use the unique sampler and evaluate on 1k games (cf. Table A1) and 10k puzzles from a different month than that used for training.
We train a vanilla decoder-only transformer without causal masking [13], with the improvements proposed in LLaMA [14] and Llama 2 [15], i.e., post-normalization and SwiGLU [49]. We use three different model configurations (with a widening factor of 4): (i) 8 heads, 8 layers, and an embedding dimension of 256, (ii) 8 heads, 8 layers, and an embedding dimension of 1024, and (iii) 8 heads, 16 layers, and an embedding dimension of 1024.
@@ A.3 Ablation Setup
We train for 5 million steps with a batch size of 1024, meaning that we train for 3.19 epochs on a smaller training dataset (M games). We use the Adam optimizer [17] with a learning rate of . We train on the dataset generated from 1 million games (cf. Table A1) for the action-value policy with 32 return buckets and a stockfish time limit of 50ms. We use the unique sampler and evaluate on K games (cf. Table A1) and K puzzles ( overlap with training set) from a different month than that used for training. We train a vanilla decoder-only transformer [13] with post-normalization, 8 heads, 8 layers, an embedding dimension of 256, a widening factor of 4, and no causal masking.
@@ A.4 Dataset Statistics
We visualize some dataset statistics in Figure A1. Note that, even though we should have the same amount of board-state data points for state-value prediction and behavioral cloning, we occasionally run into time-outs when annotating our dataset with Stockfish (due to our internal infrastructure setup) and therefore drop the corresponding record.
A.5 Playing-Strength Evaluation
Lichess
@@ A.5 Playing-Strength Evaluation
We evaluate and calibrate the playing strength of our models by playing against humans and bots on Lichess (using the lichess-bot project [50] as a basis). Our standard evaluation allows for both playing against bots and humans (see Table 1), but since humans tend to rarely play against bots the Elo ratings in this case are dominated by playing against other bots (see our discussion of how this essentially creates two different, somewhat miscalibrated, player pools in Section 4). In our case the policies in the column denoted with ‘vs. Bots’ in Table 1 have played against some humans but the number of games against humans is of total games played. To get better calibration against humans we let our largest model play exclusively against humans (by not accepting games with other bots) which leads to a significantly higher Elo ranking (see Table 1). Overall we have played the following numbers of games for the different policies shown in Table 1: 553 games for M, 169 games for M, 228 games against bots and 174 games against humans for M, 181 games for GPT-3.5-turbo-instruct, 37 games for Leela Chess Zero (policy net only), 46 games for Leela Chess Zero (value net only), 44 games for Leela Chess Zero (400 MCTS sim.), 30 games for Stockfish (50ms per move) [oracle], and 42 games for Stockfish (1.5s per board).
State-Value | Behavioral Cloning | Action-Value | |||||||
Split | Games | Records | Bytes | Records | Bytes | Records | Bytes | ||
Train | 591 897 | 43.7 MB | 589 130 | 41.1 MB | 17 373 887 | 1.4 GB | |||
5 747 753 | 422.0 MB | 5 720 672 | 397.4 MB | 167 912 926 | 13.5 GB | ||||
55 259 971 | 4.0 GB | 54 991 050 | 3.8 GB | 1 606 372 407 | 129.0 GB | ||||
530 310 443 | 38.6 GB | 527 633 465 | 36.3 GB | 15 316 914 724 | 1.2 TB | ||||
Test | 62 829 | 4.6 MB | 62 561 | 4.4 MB | 1 838 218 | 148.3 MB |
A.6 Stockfish Setup
@@
We use Stockfish 16 (the version from December 2023) throughout the paper. When we play, we use the oracle we used for training, which is an unconventional way to play with this engine: We evaluate each legal move in the position for 50ms, and return the best move based on these scores. This is not entirely equivalent to a standard thinking time of 50ms multiplied by the number of legal moves per position, as we force Stockfish to spend 50ms on moves that could be uninteresting and unexplored. We chose to keep this setup to have a comparison to the oracle we train on. Note that, when comparing the legal moves in a given position, we do not clear Stockfish’s cache between the moves. Therefore, due to the way the cache works, this biases the accuracy of Stockfish’s evaluation to be weaker for the first moves considered. Finally, due to the nature of our internal hardware setup, we use two different kinds of chips to run Stockfish: (i) to compute the Lichess Elo, we use a 6-core Intel(R) Xeon(R) W-2135 CPU @ 3.70GHz, and (ii) to compute the tournament Elo, we use a single Tensor Processing Unit (V3), as for all the other agents.
A.7 AlphaZero Setup
@@
We use the AlphaZero [7] version from 2020, with a network trained at that time [23]. We investigate three different versions: (i) policy network only, (ii) value network only and (iii) standard version with search. For (i) we do a search limited to 1 node. For (ii) we do a search limited to 1 node for every legal action and then take the . For (iii), we use the standard search from the paper, with 400 MCTS simulations (confirmed via personal correspondence with Schrittwieser et al. [23]) and the exact same UCB scaling parameters. We also take the over visit counts. Note that AlphaZero’s policy and value network have been trained on M games, whereas we trained our largest models on only M games.
A.8 Leela Chess Zero Setup
We use the latest release version (v31.0) of Leela Chess Zero [5] with the largest network officially supported at the time of submission (i.e., the ‘Large’ T82 network from https://lczero.org/play/networks/bestnets). We investigate three different versions: (i) policy network only, (ii) value network only and (iii) standard version with search. For (i) we do a search limited to 1 node. For (ii) we do a search limited to 1 node for every legal action and then take the . For (iii), we use 400 MCTS simulations. The network architecture and training protocol for T82 were not fully available; more details are now provided in Monroe and Leela Chess Zero Team [6] (written concurrently to our work) about a slightly different architecture, called the ChessFormer.
A.9 Computational Resources
@@
Our codebase is based on JAX [51] and the DeepMind JAX Ecosystem [52, 53]. We used 4 Tensor Processing Units (V5) per model for the ablation experiments. We used 128 Tensor Processing Units (V5) per model to train our large (9M, 136M and 270M) models. We used a single Tensor Processing Unit (V3) per agent for our Elo tournament.
@@ Appendix B Additional Results
B.1 Additional Ablations
We perform additional ablations using the M parameter model. As in Section 3.4, the results and conclusions drawn are used to justify our design choices and determine default configuration for model, data, and training. Table A2 shows the results.
Accuracy (%) | ||||
Ablation | Parameter | Puzzles | Actions | Kendall’s |
Stockfish limit [s] | 0.05 | 84.0 | 62.2 | 0.256 |
0.1 | 85.4 | 62.5 | 0.254 | |
0.2 | 84.3 | 62.6 | 0.259 | |
0.5 | 83.3 | 63.0 | 0.259 | |
Data sampler | Uniform | 83.3 | 63.0 | 0.259 |
Weighted | 49.9 | 48.2 | 0.192 | |
Architecture | Transformer | 83.3 | 63.0 | 0.259 |
ConvNet | 17.3 | 37.6 | 0.171 |
Stockfish time limit
We create training sets from million games annotated by Stockfish with varying time limits to manipulate the playing strength of our oracle. We report scores on the puzzle set (same for all models) and a test set created using the same time limit as the training set (different for all models). Table 2 shows that a basic time limit of seconds gives only marginally worse puzzle performance. As a compromise between computational effort and final model performance we thus choose this as our default value (for our M games dataset we need about B action-evaluation calls with Stockfish, i.e., roughly 8680 days of unparallelized Stockfish evaluation time).
Data sampler
We remove duplicate board states during the generation of the training and test sets. This increases data diversity but introduces distributional shift compared to the “natural” game distribution of boards where early board states and popular openings occur more frequently. To quantify the effect of this shift we use an alternative “weighted” data sampler that draws boards from our filtered training set according to the distribution that would occur if we had not removed duplicates. Results in Table A2 reveal that training on the natural distribution (via the weighted sampler) leads to significantly worse results compared to sampling uniformly randomly from the filtered training set (both trained models are evaluated on a filtered test set with uniform sampling, and the puzzle test set). We hypothesize that the increased performance is due to the increased data diversity seen under uniform sampling. As we train for very few epochs, the starting position and common opening positions are only seen a handful of times during training under uniform sampling, making it unlikely that strong early-game play of our models can be attributed to memorization.
Model architecture
To investigate the performance of architectures other than transformers on our benchmark, we perform a preliminary investigation with an AlphaZero-like convolutional neural network. We therefore train a residual tower with 8 layers and 256 channels (no pooling, broadcasting, or bottleneck layers), followed by a linear layer and a . We input the -by- board in a single channel and treat the other scalars in the FEN as additional channels. Table A2 shows that this particular model- and training configuration does not lead to strong results for the convolutional architecture we consider. Note that we did not extensively tune our experiment setup and would therefore expect highly domain-specific setups that more closely resemble those of AlphaZero or Leela Chess Zero to achieve much higher performance. Thus, the purpose of this experiment is to show that the vanilla transformer setup seems to achieve better performance than an “equivalently vanilla” convolutional setup.
Inference Time [ms] | ||
Agent | Average | Standard Deviation |
9M Transformer (ours) | ||
136M Transformer (ours) | ||
270M Transformer (ours) | ||
AlphaZero (policy net only) | ||
AlphaZero (value net only) | ||
AlphaZero ( MCTS sim.) | ||
Leela Chess Zero (policy net only) | ||
Leela Chess Zero (value net only) | ||
Leela Chess Zero ( MCTS sim.) | ||
Stockfish 16 (ms per move) [oracle] | ||
Stockfish 16 (s per board) |
Legal Move Accuracy (%) | |||
Prediction Target | ECO | Puzzles | Test |
Action-Value | 0.0 | 10.5 | 0.5 |
Behavioral Cloning | 100.0 | 99.6 | 99.5 |
B.2 Inference Times
Table 3(a) shows the inference times in milliseconds on 1000 random boards from the Encyclopedia of Chess Openings [22] using a single TPU (V5). Unsurprisingly, the “searchless” agents (i.e., our models and the policy networks) are orders of magnitudes faster than the agents that rely on tree search at test time. Note that the value networks are slow to evaluate since we compute every legal action individually (and sequentially) and then take the over the values.
B.3 Predicting Legal Moves
A natural question that arises when training models to play chess is whether they have actually learned the rules of the game. However, the problem of generating legal moves is tangential to our paper since we use a legal move generator (given by the environment) for our policies, as is standard in the chess AI literature (see, e.g., AlphaZero, Stockfish, or Leela Chess Zero). In fact, computing the legal move accuracy does not actually make sense for action-value prediction since the action-value function is only defined for legal moves. In contrast, for behavioral cloning the question of legal move accuracy is potentially more interesting (throughout this work, we normalize the predictive distribution to support only the legal moves except when evaluating the legal move accuracy) since the model is trained to output the oracle (and therefore a valid) action.
We investigate the legal move accuracy for action-value and behavioral cloning in Table 3(b) (the legal move accuracy is undefined for state-value prediction since we can only evaluate boards reachable via legal moves). Unsurprisingly, Table 3(b) shows that the action-value predictor has a very low legal move accuracy, since its outputs for illegal moves are undefined (i.e., never adjusted during training). In contrast, our behavioral cloning policy has learned to almost always predict a legal move, even for highly out-of-distribution puzzle boards, showing that our networks have the capacity to learn the rules given our training protocol and dataset – if explicitly trained to do so (i.e., unlike action-value prediction).
B.4 Loss Curves
@@
In Figure A3 we visualize the train and test loss curves for the scaling experiment from Section 3.3. In line with the results shown in the main paper we observe that on the smallest training set, models with M parameters start to overfit but not for the larger training sets. Except for the overfitting cases we observe that larger models improve both the training and test loss, regardless of training set size, and that larger training set size improves the test loss when keeping the model size constant.
B.5 Predictor-Target Comparison
@@
Throughout all these results we observe lower performance of the behavioral cloning policy, despite being trained on a comparable number of data points as the state-value policy. The main hypothesis for this is that the amount of information in the behavioral cloning dataset is lower than the state-value dataset, since we throw away any information in the state- or action-values beyond the index of the oracle action. We suspect that training on the full action distribution of the oracle (with cross-entropy loss), rather than the best action only would largely close this gap, but we consider this question beyond the scope of this paper and limit ourselves to simply reporting the observed effect in our setting.
B.6 Fischer Random Chess
As a measure of out-of-distribution generalization, we evaluate the performance of our 9M transformer on Fischer random chess333https://en.wikipedia.org/wiki/Fischer_random_chess (Chess960) via Lichess (using the same setup as for Table 1). Fischer is a variation of the game of chess, which employs the same board and pieces as classical chess, but the starting position of the pieces on the players’ home ranks is randomized, following certain rules. The random setup makes gaining an advantage through the memorization of openings impracticable; players instead must rely more on their skill and creativity over the board. Thus, Fischer random chess can be viewed as a measure of whether our models have learned general tactics (for the early stages of the game). While our bot attains a blitz Elo of 2054 (cf. Table 1), it only manages to play Fischer random chess at 1539 Elo (against bots), suggesting that the distribution shift is too high for our models to handle.
B.7 Tactics
@@
B.8 Playing Style
We recruited chess players of National Master level and above to analyze our agent’s games against bots and humans on the Lichess platform. They made the following qualitative assessments of its playing style and highlighted specific examples (see Figure A7).
@@ B.8 Playing Style
The agent has an aggressive enterprising style where it frequently sacrifices material for long-term strategic gain. The agent plays optimistically: it prefers moves that give opponents difficult decisions to make even if they are not always objectively correct. It values king safety highly in that it only reluctantly exposes its own king to danger but also frequently sacrifices material and time to expose the opponent’s king. For example 17 .. Bg5 in game B.8.1 encouraged its opponent to weaken their king position. Its style incorporates strategic motifs employed by the most recent neural engines [7, 54]. For example it pushes wing pawns in the middlegame when conditions permit (see game B.8.2). In game B.8.3 the agent executes a correct long-term exchange sacrifice. In game B.8.4 the bot uses a motif of a pin on the back rank to justify a pawn sacrifice for long term pressure. Game B.8.5 features a piece sacrifice to expose its opponent’s king. The sacrifice is not justified according to Stockfish although the opponent does not manage to tread the fine line to a permanent advantage and blunders six moves later with Bg7.
@@
The agent has a distinct playing style to Stockfish: one analyzer commented “it feels more enjoyable than playing a normal engine”, “as if you are not just hopelessly crushed”. Indeed it does frequently agree with Stockfish’s move choices suggesting that the agent’s action-value predictions match Stockfish’s. However the disagreements can be telling: the piece sacrifice in the preceding paragraph is such an example. Also, game B.8.6 is interesting because the agent makes moves that Stockfish strongly disagrees with. In particular the agent strongly favours 18 .. Rxb4 and believes black is better, in contrast Stockfish believes white is better and prefers Nd4. Subsequent analysis by the masters suggests Stockfish is objectively correct in this instance. Indeed on the very next move the agent has reversed its opinion and agrees with Stockfish.
@@
The agent’s aggressive style is highly successful against human opponents and achieves a grandmaster-level Lichess blitz Elo of 2895. However, against other engines its estimated Elo is far lower, i.e., 2299. Its aggressive playing style does not work as well against engines that are adept at tactical calculations, particularly when there is a tactical refutation to a sub-optimal move. Most losses against bots can be explained by just one tactical blunder in the game that the opponent refutes. For example Bxh3 in game B.8.7 loses a piece to g4.
@@
Finally, the recruited chess masters commented that the agent’s style makes it very useful for opening repertoire preparation. It is no longer feasible to surprise human opponents with opening novelties as all the best moves have been heavily over-analyzed. Modern opening preparation amongst professional chess players now focuses on discovering sub-optimal moves that pose difficult problems for opponents. This aligns extremely well with the agent’s aggressive, enterprising playing style which does not always respect objective evaluations of positions.
B.8.1 King weakening game